The sum of 2440 consists of two numbers 100 and 10. If 100 is six times that of 10. How many 100 are there?

1
Expert's answer
2022-06-08T01:36:04-0400

Let the amount of number 10 is "x", then amount of number 100 is "6x." We get

"2440=x*10+6x*100\\
2440=610x\\
x=4"

Answer: the number 100 is 6*4=24 times.